Home Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  Alibava/Mapsa testbeam at DESY in April 2017 (Axel, Grossi)  Not logged in ELOG logo
Message ID: 98     Entry time: Mon Apr 17 10:14:15 2017
Author: Axel Koenig 
Project: common 
Subject: Run summary Baby 
Run Number:  
Events:  
XYPosition:  
Cooling: No 
Data: Unknown 
Alibava run Telescope run File existence check Run type Triggers Remarks
001 562 yes first run 250000 none
002 563 yes central position shifted by -1 cm vertically 250000 none
003 564 yes central position shifted by + 1cm 250000 none
004 565 yes central position angle +10 deg 250000 none
005 568 yes central position angle -10 deg 250000 none
006 570 yes unknown (maybe long run over night that crashed) ? data may be corrupted
007 579 yes bias scan -10V 450000 run crashed data only valid up to 274915
008 583 yes bias scan -20V 175000 none
009 586 yes bias scan -30V 135000 none
010 587 yes bias scan -40V 135000 none
011 588 yes bias scan -50V 135000 none
012 589 yes bias scan -60V 135000 none
013 590 yes bias scan -65V 112000 none
014 591 yes bias scan -70V 112000 none
015 592 yes long run over night 2500000 none
016 593 yes bias scan -75V 112000 none
017 594 yes bias scan -80V 112000 none
018 595 yes bias scan -90V 112000 none
019 596 yes bias scan -110V 112000 none
020 597 yes bias scan -120V 112000 none
021 598 yes bias scan -130V 112000 none
022 599 yes bias scan -140V 112000 none
023 600 yes bias scan -150V 112000 none

Central beam position: x = 102 mm, y = 22.5 mm, rot = 240 deg

Current was stable for all runs i.e. 500 +- 10 nA at 100 V

Trigger efficiency ~ 90% for all runs (except bias scan).

Temperature in area 20+-1 deg.

Telescope chiller off for all runs.

ELOG V3.1.4-966e3dd